OpenStudy (anonymous):

The three Math Idol judges have been eliminating contestants all day! The number of one-step equations and two-step equations who have been eliminated today is equal to 840! If three times the number of one-step equations minus twice the number of two-step equations is equal to 870, how many one-step equations auditioned today?

OpenStudy (lalaly):

x+y=840 3y-2x=870

myininaya (myininaya):

one step equations=o two step equations=t we have o+t=840 and 3o-2t=870

OpenStudy (anonymous):

90? "/